OpenStudy (anonymous):

Which equation is the result of solving 9 + 3x = 2y for x?

OpenStudy (anonymous):

9 + 3x = 2y 3x = 2y - 9 x = 2/3y - 3 You're ultimate goal is to get x on its own. In the first step I subtracted the 9 from the left hand side and because I did that I MUST do it to the right hand side as well. If you don't the equation becomes unbalanced. In the second step, because the x was multiplied by 3 I had to do the opposite to get the x on it's own and because I divided the left by 3 I must also divide the right by 3. The main idea is to do the opposite until you get what you want on its own and for every step you take you must do the same thing to the other side of the equation to balance it.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No proplem :)